### v2

# Old Quantum Mechanics by Bohr and Sommerfeld from a Modern Perspective

Abstract.
We review Bohr’s atomic model and its extension by Sommerfeld from a mathematical perspective of wave mechanics. The derivation of quantization rules and energy levels is revisited using semiclassical methods. Sommerfeld-type integrals are evaluated by elementary techniques, and connections with the Schrödinger and Dirac equations are established. Historical developments and key transitions from classical to quantum theory are discussed to clarify the structure and significance of the old quantum mechanics.
Key words and phrases:
Bohr model, Bohr–Sommerfeld quantization rule, Schrödinger equation, Dirac equation, WKB method.1991 Mathematics Subject Classification:
Primary 81-01, 81-03. Secondary 81CDedicated to the 100th anniversary of the birth of wave mechanics [52].
1. Introduction
The study of blackbody radiation and the quantum theory that arose from it laid the foundation for Bohr’s atomic model, which became a major step in understanding the structure of the atom about a century ago. By recognizing the quantum nature of energy and the discrete energy levels of electrons, Planck [47], Einstein [18], Rutherford [50], and Bohr [7] began to explain the behavior of light and matter at the atomic level, thereby paving the way for the development of quantum mechanics.
Among the key sources on the so-called “Old Quantum Mechanics” of Bohr and Sommerfeld are the classic publications [1], [8], [20], [30], [31], [40], [44], [49], [56], [57], [66], and several educational videos [9]. In this paper, we aim to explore the following topics from a mathematical perspective:
Topics: (i) The Bohr model: circular orbits of electrons in hydrogen-like atoms; derivation of the Bohr formula (Nobel Prize in Physics, 1922 [7]). (ii) Wilson and Sommerfeld: quantization rules for multi-dimensional periodic systems via classical action; Sommerfeld’s relativistic formula for elliptical orbits. (iii) Elementary evaluation of Sommerfeld-type integrals. (iv) Additional examples and a resolution of the “Sommerfeld puzzle”. (v) Appendix A: Vector calculus tools for uniform circular motion. (vi) Appendix B: Instability of the hydrogen atom in classical physics due to the electron’s fall into the center, as predicted by Rutherford’s model. (vii) Appendix C: Independent evaluation of Sommerfeld-type integrals using parameter differentiation.
Brief history: The fine structure of hydrogen atom spectral lines was discovered by Albert A. Michelson in 1887 [42, 43]. When his ether-wind experiments failed, he turned to spectroscopy and found that the prominent line of the Balmer series was actually a doublet [6, 49]. The electron was discovered by J. J. Thomson in 1897 [63]. Rutherford proposed the planetary model of the atom in 1911. In 1916, Arnold Sommerfeld extended the quantization rules of the ‘old’ quantum theory to the relativistic hydrogen atom [57] (see also [23, 29, 44, 49]). An exact solution was achieved only in 1928 by C. G. Darwin [11] and W. Gordon [22] after the discovery of the Dirac equation [14, 16] – astonishingly, the new result precisely matched the ‘old’ Sommerfeld formula – the so-called “Sommerfeld Puzzle” [6]!
Traditional physics textbooks skip the semiclassical derivation of the Sommerfeld fine structure formula due to its complexity and because an accurate and elegant solution is only available through relativistic quantum mechanics. The semiclassical approach, while providing a simplified model, requires a careful and sometimes challenging analysis, making it less accessible for introductory physics courses.
2. Bohr’s Atomic Model
The second Newton law for the uniform circular motion of a charged particle, such as an electron in the static Coulomb field of a heavy ion, with the positive charge states
by (7.1) from Appendix 7. (Here and are the electron mass and the absolute value of its electric charge in the centimeter-gram-second (cgs) system units, respectively.) For the electron linear momentum, one gets
and the total energy is given by
which is exactly one half of the potential energy by the virial theorem.
Niels Bohr [1, 8, 31], following experiments of Nickolson [39], suggested to quantize the corresponding electron angular momentum
in terms of the reduced Planck constant in the cgs units. For a uniform circular motion, vectors and are perpendicular to each other (7.3) (see Figure 1). As a result, he obtained the so-called Bohr’s orbits:111In terms of the de Broglie wavelength the quantization rule states that the length of the orbit equals where .
and the corresponding electron discrete energy levels:
where is the principal quantum number.
resulting in (2.5). In a similar fashion, by (2.3) and (2.5):
which completes the derivation of Bohr’s discrete energy formula (2.6).
In Bohr’s atomic model, electrons in the orbits (2.5) are stable and do not radiate energy. (An instability of the original Rutherford atom is discussed in Appendix 8.) Electrons can transition between energy levels (2.6) by absorbing or emitting photons (light quanta) with specific energies. The energy of the photon corresponds to the difference in energy between the initial and final levels [31] (see, for example, Figure 2).
Mendeleev’s Periodic Table and the Bohr model are two significant developments in understanding the structure of elements and atoms — Mendeleev’s table, created in 1869, organized elements by increasing atomic weight and recurring chemical properties [56, pp. 2–3]. The Bohr model, proposed in 1913 [8], provided a model for the structure of the atom, depicting electrons orbiting the nucleus in fixed energy levels. The simplified way of understanding the electronic structure of atoms is directly related to the arrangement of elements in the periodic table.
3. Wilson and Sommerfeld Quantization Rules in Wave Mechanics
Topics to review: The spherical harmonics [35, 46, 64], Schrödinger equation [13, 35, 51], relativistic Schrödinger and Dirac equations [3, 13, 51, 62]; the spinor spherical harmonical harmonics [61, 62, 64] and separation of variables for the Dirac equation in a central field, semiclassical approximation [46, 51].
Brief history: As an extension of Bohr’s rules, Wilson [66] and Sommerfeld [57] independently proposed a method of quantizing action integrals in classical mechanics for a multi-dimensional periodic system over one period of motion.
In the development of quantum theory, the Bohr–(Wilson)–Sommerfeld quantization rule served as an original “bridge” between classical and quantum mechanics (for historical details, see [39], [40], [44], [49], [57], [58], and [66]). Nowadays, we use the Schrödinger [55] and Dirac [15] wave equations for the corresponding Kepler problems. How did Schrödinger derive his celebrated equation and subsequently apply it to the hydrogen atom? According to his own testimony [52, 53, 54], de Broglie’s seminal work on a wave theory of matter (1923–24) [12] and Einstein’s work on ideal Bose gases (1924–25) laid the foundation for the discovery of wave mechanics (see also [3] and [41]).
The phenomenological quantization rules of ‘old’ quantum mechanics [57, 66] are derived in modern physics from the corresponding wave equations via the so-called semiclassical approximation [4], [10], [32], [35], [46], and [65] – the Wentzel–Kramers–Brillouin (WKB) method.
This approximation refines the Bohr–Sommerfeld quantization rule within wave mechanics. The WKB method, which provides approximate solutions to wave equations, leads to a quantization condition similar to the Bohr–Sommerfeld rule, with a slight modification involving a phase shift. During separation of variables, the quantization of angular momentum and spin is exact, since the concept of spin is already embedded into the structure of the corresponding wave equation [33].
Let us recall the one-dimensional stationary Schrödinger equation:
For a particle in a central field, the corresponding -wave equations can be separated in spherical coordinates. Then one usually obtains a radial equation of the form:
where is continuous along with its first and second derivatives for . These equations can be approximately solved by the WKB method.
The corresponding wave functions, their relations to Airy functions [17], quantization rules, and further technical details are discussed in [46, 51], and elsewhere. (We recommend the reader to review sections 19, pp. 235–251, and 28, pp. 178–188, on the semiclassical approximation of [46, 51], respectively.)
It is well known that the traditional semiclassical approximation breaks down near for central fields. However, using the change of variables and transforms the equation into the new form:
where
This is known as Langer’s transformation [36, 37]. As (i.e., ), the function varies slowly near the constant:
Thus, and its derivatives vary slowly for large negative .
The WKB method can be applied to the new equation, and in the original equation one replaces with:
The Bohr–Sommerfeld quantization rule, derived for example in [46] and [51], takes the form:
provided .
For all Coulomb problems under consideration, we utilize a generic integral, originally evaluated by Sommerfeld [58] using complex analysis: If
then:
with . (See Section 4 and Appendix 9 for two independent elementary evaluations of this integral.)
As a result, for the discrete energy levels, we obtain the following generic equation:
which is valid for all Coulomb-type problems under consideration and more [59, 60].
Examples. For the well-known case of the non-relativistic Coulomb problem, one obtains the following equation in dimensionless units [35, 51]:
where is the quantized orbital angular momentum.
In the Bohr–Sommerfeld quantization rule, one must take:
Identifying parameters in the generic integral (3.7), we have:
Substituting into the quantization rule (3.9), we find:
Solving for , this yields the exact energy levels for the non-relativistic hydrogen-like problem:
Here, is the principal quantum number, recovering Bohr’s formula for discrete energy levels as given in Eq. (2.6) in the semiclassical approximation.
Our main goal is to analyze the corresponding relativistic problems. In the case of the relativistic Schrödinger equation, one writes [3]:
(see Figure 3 for the original version222Schrödinger’s notebooks are reproduced in the Archive for the History of Quantum Physics (AHQP); for more details, see [27, 41]. ) and applies Langer’s transformation to define the effective momentum:
Here we identify the parameters:
Applying the Bohr–Sommerfeld quantization condition (3.9) yields:
which gives the exact relativistic energy levels:
with
In the non-relativistic limit (or ), one obtains [13, 51]:
where is the corresponding non-relativistic principal quantum number (see also [2] for a complementary Mathematica notebook).
In this Taylor’s expansion: – The first term corresponds to the rest energy . – The second term gives the non-relativistic Schrödinger energy eigenvalue. – The third term represents the so-called fine structure, lifting degeneracy between states with the same but different .
However, Schrödinger’s relativistic approach fails to accurately describe the fine structure of hydrogen-like atoms (hydrogen, ionized helium, doubly-ionized lithium, etc.). For instance, the total fine-structure splitting at is a factor of too large compared to Sommerfeld’s prediction, which agrees with experimental observations.333The maximum spread in the fine-structure levels occurs for and with total angular momentum and in Eqs. (3.19) and (3.24), respectively [13, 51]. The ratio of these spreads is:
Sommerfeld’s fine structure formula for the relativistic Coulomb problem represents one of the most significant achievements of the ‘old’ quantum mechanics. Here, we derive this result in the semiclassical approximation using the radial Dirac equations (separation of variables in spherical coordinates is discussed in detail in Refs. [5], [46], [61], and [62]).
In dimensionless units, one of the second-order differential equations for the Dirac spinor component takes the form:
while the second equation can be obtained by the substitution (see Eqs. (3.81)–(3.82) in Ref. [62]).
Applying Langer’s transformation leads to an effective momentum function:
Thus, for the Dirac equation we identify:
Applying the Bohr–Sommerfeld quantization rule (3.9) , one arrives at (3.16) and the corresponding energy spectrum occurs:
with the adjustment as discussed in [46, 62]. Here, and in Dirac theory,
where is the total angular momentum (including spin).
In the non-relativistic limit (), Dirac’s formula yields [5, 13, 51, 62]:
where is the principal quantum number for the hydrogen-like atom (see also [2]).
In this expansion: – The first term is the rest mass energy of the electron, . – The second term coincides with the non-relativistic Schrödinger energy. – The third term gives the fine structure correction, originating from the spin-orbit interaction in the Pauli approximation.
This prediction agrees with experimental data on fine-structure splittings for the hydrogen-like systems.
Hence, the Sommerfeld fine structure formula can nowadays be derived semiclassically from the radial Dirac equations. This derivation elucidates the quantization rules of Bohr and Sommerfeld [58, 59], introduced a decade before the concept of spin444The concept of electron spin was introduced by G. E. Uhlenbeck and S. Goudsmit in a letter published in Die Naturwissenschaften; the issue of 20 November 1925; see [38] for more details. , and bind them with modern quantum theory. Indeed, the classical relativistic Hamiltonian does not include a spinning electron, which creates an ambiguity in the Bohr–Sommerfeld quantization (see [23, 48]).
For a full analytical solution, including the non-relativistic limit, see Refs. [19, 46, 61, 62] (based on the Nikiforov–Uvarov method), as well as standard texts [13, 51].
4. Evaluation of the Sommerfeld-Type Integrals
For all problems under consideration, we utilize the generic integral (3.7)–(3.8), originally evaluated by Sommerfeld using complex integration [58] (see Figure 4). We now discuss an elementary evaluation of this integral [3]. Integrating by parts on the left-hand side of (3.8), one obtains:
For the penultimate integral, we write:
Next, applying the substitution in the final integral of (4.1), we obtain:
where we completed the square and evaluated a standard definite integral. (Alternatively, one may interchange and and follow a similar route.)
Combining the results from the last two integrals completes the proof.
5. Further Examples and Resolution of “Sommerfeld’s Puzzle”
As is well known, Bohr introduced his semiclassical quantization rules for hydrogen-like atoms based on classical circular motion, but Sommerfeld extended these ideas to relativistic elliptical orbits. Measurements of the fine structure by Paschen were interpreted as experimental tests of the special theory of relativity [6, 29]. The exact solution was obtained, for the first time, by C. G. Darwin [11] and W. Gordon [22], only after the discovery of the Dirac equation [14, 16] – the new answer was precisely the ‘old’ Sommerfeld formula (3.22)!
Werner Heisenberg [24] called this agreement a ‘miracle’ and wrote:“It would be intriguing to explore whether this is about a miracle or it is the group-theoretical approach which leads to this formula”[25]. Erwin Schrödinger, in a letter from 1956, commented:“This is a fortuitous coincidence”[67]. As shown in [60], Schrödinger appears to be right. The “Sommerfeld Puzzle” [6] has been resolved and extended to a certain class of multi-dimensional problems with different symmetry groups.
Indeed, the generic (WKB-based) rule (3.9) is also valid for the exact energy levels obtained via the Nikiforov–Uvarov approach [19, 60] for all Coulomb problems under consideration. Other examples include quantum harmonic oscillators, and systems with Kratzer and Pöschl–Teller potentials [19]. (See also [34] for an extension of Schrödinger’s coherent states [53].)
In connection with Sommerfeld’s fine-structure formula Erwin Schrödinger testified, inter alia, in the same letter dated 29th February, 1956 [67]:“… you are naturally aware of the fact that Sommerfeld derivation of the fine-structure formula provides only fortuitously the result demanded by the experiment. One may notice then from this particular example that newer form of quantum theory (i.e., quantum mechanics) is by no means such an inventible continuation of the older theory as is commonly supposed. Admittedly the Schrödinger theory, relativistically framed (without spin), gives a formal expression of the fine-structure formula of Sommerfeld, but it is incorrect owing to the appearance of half-integers instead of integers. My paper in which this is shown has … never been published; it was withdrawn by me and replaced by non-relativistic treatment… The computation [by the relativistic method] is far too little known. It shows in one respect how necessary Dirac’s improvement was, and on the other hand it is wrong to assume that the older form of quantum theory is ‘broadly’ in accordance with the newer form.”
Note. It can be summarized that only after the “Two Quantum Revolutions” the ambiguity of quantization of the Kepler problem in the ‘old quantum mechanics’ [23, 48] was resolved, when the spherical symmetry had been taking into account by separation of variables and the corresponding radial equations were derived, without any assumptions, subject to further WKB approximation under the Langer transformation. Schrödinger seems to have done it first in 1925 in his unpublished notes but he used equation (3.14), which corresponds to the spin zero particle (see [3], Figure 3 for his original notebook, and [33] for further details on the concept of spin and the wave equations). As was later attested, due to discrepancy with experimental data, he never published this work (see also [3], Appendix D, for his letter to Weyl).
6. Conclusion
Traditional textbooks [13, 35, 51] do not discuss the derivation of the Sommerfeld fine structure formula in the semiclassical approximation, and now the reader can understand why! Indeed, de facto, there are three different levels of complexity, as in the historical “Three Quantum Revolutions”: – Elementary one, in the Bohr model of atom. – Introductory quantum mechanics for the non-relativistic hydrogen atom. – And finally, relativistic quantum theory, the Dirac equation, for the fine structure formula. The WKB method can be applied only after a thorough study of basic properties of the Dirac equation, including the construction of spinor spherical harmonics [61, 62, 64] followed by a non-trivial separation of variables in spherical coordinates. Our notes, while not perfect, may help the reader fill in these gaps. The resolution of the “Sommerfeld Puzzle” is also of interest [60]. Needless to say, practical use of computer algebra [2, 19] will definitely help students in learning quantum physics, for instance, when performing tedious calculations.
7. Appendix A: Velocity, Acceleration and Angular Momentum for
the Uniform Circular Motion
For a uniform circular motion one gets
8. Appendix B: Instability of a Hydrogen Atom in Classical Physics
As is known [9, 49, 56]a rotating electron in Rutherford’s planetary model must fall into the nucleus, the spiral-in, according to the laws of classical mechanics and electrodynamics. The total instantaneous power emitted over all solid angles is given by the well-known result of Larmor [26]:
where cm/s is the speed of light in cgs units. Here, according to (2.1)–(2.3),
and thus,
On the other hand, from the virial theorem (2.3), one gets:
Equating (8.3) and (8.4), we obtain:
Integrating both sides yields:
where cm is the first Bohr radius of a hydrogen atom as given in (2.5). Therefore, an electron in Rutherford’s model will fall into the nucleus in less than a nanosecond:
The electron velocity on the first Bohr orbit can be estimated as follows
(a non-relativistic motion) and, for the time of one revolution, we arrive at
Therefore, the total number of rotations before falling into the center can be estimated as
The spiral-in time, is much longer that the orbital time, so treating the spiral-in as a succession of circular orbits in a hydrogen atom is plausible (see [49, 56] for more details) .
9. Appendix C: An Independent Evaluation of the Sommerfeld-type Integrals
On the contrary, one can use the technique of differentiation with respect to parameters for the familiar integrals related to the Bohr–Sommerfeld quantization rule [58, 59]. As is well known, if
then
In the WKB case, the last two terms vanish because the limits are turning points where the integrand vanishes [21].
We now apply this procedure for an independent evaluation of the “Sommerfeld-type” integrals discussed in this note [60]. Indeed,
provided , one finds:
As a result,
and by integration,
Acknowledgments. We are grateful to Dr. Sergey I. Kryuchkov, Dr. Eugene Stepanov, and Dr. Andreas Ruffing for their help and valuable comments.